Consider writing onto a computer disk and then sending it through a certifier that counts the number of missing pulses. Suppose this number X has a Poisson distribution with parameter ? = 0.6. (Round your answers to three decimal places.)

 (a) What is the probability that a disk has exactly one missing pulse?  

 (b) What is the probability that a disk has at least two missing pulses? 

 (c) If two disks are independently selected, what is the probability that neither contains a missing pulse?   


 You may need to use the appropriate table in the Appendix of Tables to answer this question.
Answer #1
a) P(X = 1) = .6^1 * e^-.6 / 1! = .3292 (b) P(at least 2) = 1 - P(X = 0) - P(X = 1) = 1 - [.6^0 * e^-.6 / 0!] - .3292 = 1 - .5488 - .3292 = .122 (c) P(X = 0) * P(X = 0) = .5488^2 = .30118
Answer #2
P (X = k) = (e^-lambda * lambda^k )/ fact(k) a) P(X =1) = e^-0.6 * (0.6) = 0.329 b)P(X>=2) = 1 - P(X = 0) - P(X =1) = 1 - 0.549 - 0.329 = 0.122 c) P(X =0) * P(Y =0) = (0.549) * (0.549) = 0.301